Sample Title

This is a short example of a complete Texinfo file, version 1.0. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2003 Free Software Foundation, Inc.

Table of Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 First Chapter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Appendix A Unified index . . . . . . . . . . . . . . . . . . . 3 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

Preface 1

Preface 1 First Chapter

This is the first chapter.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover texts being "A GNU Manual," and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled "GNU Free Documentation License."

(a) The FSF’s Back-Cover Text is: "You have freedom to copy and modify this GNU Manual, like GNU software. Copies published by the Free Software Foundation raise funds for GNU development." The first part of this master menu lists the major nodes in this Info document, including the command and concept indices. The rest of the menu lists all the lower level nodes in the document. Here is a numbered list.

1.

This is the first item. 2.

This is the second item.

Appendix A: Unified index 3

Appendix A Unified index

Index 4

Index

C

chapter, first . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

I

index entry, another. . . . . . . . . . . . . . . . . . . . . . . . . . . 2